# Bootstrapping World Models from Dynamics Models in Multimodal Foundation Models

### v1

@@ Abstract
To what extent do vision-and-language foundation models possess a realistic world model (observation $\times$ action $\rightarrow$ observation) and a dynamics model (observation $\times$ observation $\rightarrow$ action), when actions are expressed through language? While open-source foundation models struggle with both, we find that fine-tuning them to acquire a dynamics model through supervision is significantly easier than acquiring a world model. In turn, dynamics models can be used to bootstrap world models through two main strategies: 1) weakly supervised learning from synthetic data and 2) inference time verification. Firstly, the dynamics model can annotate actions for unlabelled pairs of video frame observations to expand the training data. We further propose a new objective, where image tokens in observation pairs are weighted by their importance, as predicted by a recognition model. Secondly, the dynamics models can assign rewards to multiple samples of the world model to score them, effectively guiding search at inference time. We evaluate the world models resulting from both strategies through the task of action-centric image editing on Aurora-Bench. Our best model achieves a performance competitive with state-of-the-art image editing models, improving on them by a margin of $15\%$ on real-world subsets according to GPT4o-as-judge, and achieving the best average human evaluation across all subsets of Aurora-Bench [^1] Our code and models are available at https://github.com/yfqiu-nlp/vlm-world-model..
@@ 1 Introduction
World models (observation $\times$ action $\rightarrow$ observation) (2018_ha_schmidhuber_recurrent_world_model, ; agarwal2025cosmos, ; bruce2024genie, ; brooks2024videoGenerationWorldSimulator, ) can be successfully trained to simulate future trajectories given the history of past observations and actions. World models are instrumental in training embodied agents to endow them with specific abilities (qin2024worldsimbench, ), such as grounding on affordances (brohan2023can, ), spatio-temporal reasoning (huanginner, ; li2025imagine, ), and planning (reed2022generalist, ; yang2023learning, ; Hafner2025, ). However, learning a specialised world model is challenging. Firstly, it requires a large amount of real-world data (liu2024world, ) and even this data volume may be insufficient within the confines of the current training paradigm (motamed2025DoGenerativeModelsLearn, ). Secondly, the benefit of creating a separate world model to train a downstream embodied agent remains unclear because of possible compounding errors between the two models. Conversely, foundation models, such as vision-language models (VLMs), are already imbued with plenty of real-world knowledge of both action (in language form) and perception (in vision form), because of their large-scale pre-training. While such knowledge is not straightforward to elicit (gao2024physicallyGroundedVLM, ; qiu2024temporallyGrounded, ; abdou-etal-2021-language, ), we propose investigating a promising alternative to specialised world models, by enhancing the knowledge implicitly stored inside foundation models.
@@
Firstly, we probe whether native VLMs already contain reliable world models, facilitated by model designs that combine various modalities into a unified representation, i.e., sequences of tokens (team2024chameleon, ). In particular, we frame the assessment of world models as the ability to solve action-centric image editing tasks (krojer2024aurora, ). In such tasks, the model predicts the next observation given the previous observation and an action expressed as a language instruction. Based on our evaluation, we empirically demonstrate that existing open-source models do not prefer ground-truth trajectories compared to adversarially generated ones. Hence, we verify that the world model implicit in the original VLMs per se is not well grounded on real-world trajectories (gao2024physicallyGroundedVLM, ; qiu2024temporallyGrounded, ; abdou-etal-2021-language, ).
Figure 1: Illustration of our two strategies to bootstrap a world model from a dynamics model in Vision-Language Models: (i) synthesising trajectories for weak supervision (left) and (ii) inference-time verification of candidate observations (right).
@@
Surprisingly, we also find that acquiring a dynamics model (observation $\times$ observation $\rightarrow$ action) via supervised fine-tuning is substantially easier than directly acquiring a world model (observation $\times$ action $\rightarrow$ observation). Inspired by this observation, we propose two strategies to bootstrap the world model from the dynamics model in a given VLM, namely (i) learning from synthetic trajectories in videos automatically labelled with actions by the dynamics model; and (ii) test-time verification of predicted observations sampled from the world model through the dynamics model.
@@
For the weak supervision strategy, which is reminiscent of (baker2022videoPretraining-vpt, ), we use a dynamics model fine-tuned on the Aurora dataset krojer2024aurora to annotate motion key-frames pairs extracted from real-world videos with actions (in language form). Around 45 hours of unlabelled videos are sourced from movements-in-time (monfortmoments-MIT, ), Kinetics700 (kinetics700, ; carreira2019kinetics700Note, ) and UCF-101 (soomro2012ucf101, ). Together with the ground-truth trajectories in Aurora, the synthesised trajectory triplets (observation $\times$ annotated action $\rightarrow$ observation) are then used for supervised fine-tuning of the VLM world model. To effectively train the world model, we additionally propose a loss-weighting method which weights the loss of each image token according to the visual difference between the ground-truth source and target observations, as estimated by a recognition model. In the verification strategy, we show how using the VLM dynamics model to assign rewards to multiple samples generated by the VLM world model can effectively guide search at inference time.
@@
We conduct an extensive evaluation on MagicBrush, Action-Genome, Something-Something, WhatsUp and Kubric in Aurora-Bench (krojer2024aurora, ). We focus on Chameleon-7B as the best available open-source foundation model, and transform it into a world model (CWM; Chameleon World Model). We show that thanks to the synthetic data strategy to bootstrap world models from dynamics models, our general-purpose CWM can achieve an overall performance superior to state-of-the-art diffusion models specialised for image editing. In particular, CWM improves GPT4o-as-a-judge scores on the Something-Something, Action-Genome, and Kubric subsets of Aurora by 15%, 15% and 7%, respectively. Similarly, human evaluators rate CMW image editing consistently better. Inference-time verification can also improve Aurora-finetuned Chameleon to a comparable degree as data synthesis, providing an effective training-free bootstrapping method. In some cases, it can even be combined with trajectory synthesis for compounded gains.
To summarise our contributions:
  • We empirically show that VLMs like Chameleon-7B do not exhibit a clear preference for ground-truth real-world trajectories over heuristic-generated incorrect ones.
  • We propose two strategies to bootstrap a world model from a dynamics model inside VLMs: (i) learning from unlabelled videos annotated with actions by a dynamics model, and (ii) verifying the generated observations with the dynamics model at inference time.
  • We conduct extensive evaluations on Aurora-Bench: both GPT4o-as-a-judge and human raters demonstrate the effectiveness of our methods with a considerable margin compared to the state-of-the-art image editing models.
Figure 2: Comparison of predicted negative log-likelihoods (lower values indicate stronger model preference) for ground-truth real-world trajectories versus four types of negative trajectories. Top: Action prediction task for the dynamics model (observation $\times$ observation $\rightarrow$ action). Bottom: Next observation prediction task for the world model (observation $\times$ action $\rightarrow$ observation). The legend shows the percentage of times the model prefers the ground-truth trajectory ($\uparrow$) over the negatives ($\downarrow$).
@@
2 VLMs Lack a Consistent Preference for Real-World Trajectories
@@
The first research question we investigate in this paper is: To what extent do VLMs exhibit a preference for token sequences of actions and observations that align with real-world trajectories?
@@
To address this question, we evaluate a VLM on ground-truth trajectories from 5 subsets of Aurora-Bench (krojer2024aurora, ): MagicBrush, Something-Something, Action-Genome, Whatsup, and Kubric. Each subset contains 50 trajectory triplets of the form ($o_{s},a,o_{t}$), where $o_{s}$ is the source observation, $a$ the action text, and $o_{t}$ the next observation. Specifically, we test Chameleon-7B (team2024chameleon, ) as a VLM. [^2] In our preliminary experiments, where we compared Chameleon-7B with alternative VLMs such as Vila-U (wu2024vilaU, ) (see Table 1), it emerged as the best open-source VLM that allows for interleaving text and images.
@@
We then manually curate four types of negative trajectories using heuristic rules: two that manipulate the observation component of the trajectory triplet, and two that manipulate the action component. We design two kinds of action-level manipulation: 1) Random Action: for a given pair of observations, we substitute the original action with another randomly sampled within the same subset. 2) Counterfactual Action (CF): through a Large Language Model (LLM), [^3] Specifically, Qwen-2-7b-Instruct. we rewrite the action to contradict the original one. We also test the following observation-level manipulations. 3) Copy Observation: we directly copy the source observation as the target observation. 4) Inverse Observation: we swap the source and target observations.
@@
If the VLM were well aligned with ground-truth trajectories, it should display a clear preference by assigning them a higher likelihood than the negative alternatives. In Figure 2, we compare the negative log-likelihood Chameleon-7B (team2024chameleon, ) assigns to each ground-truth trajectory against its corresponding manipulated one. We evaluate the VLM in two tasks: action prediction (i.e., as a dynamics model) and next-observation prediction (i.e., as a world model). For each kind of negative trajectory, we also indicate the percentage of samples where the model favours the reference trajectory ($\uparrow$ Ref) versus the negative trajectory ($\downarrow$ Type). Additionally, we report the Pearson correlation coefficient between the predicted likelihoods for the reference and negative trajectories; a higher correlation suggests that the model struggles to distinguish between the two, assigning them similar likelihoods.
@@
From Figure 2, it emerges that Chameleon-7B displays a very limited preference for the ground-truth trajectories in a zero-shot setting. In the action prediction task (top panel), there is a slightly higher tendency to favour the ground-truth; however, even in the best case (counterfactual action), the model prefers the reference in only 58.1% of the samples. The high correlation in likelihoods indicates that the VLM struggles also on visual manipulations. In the next-observation prediction task (bottom panel), the VLM mostly fails in effectively differentiating the ground truth from the negatives. An exception to this is the copy manipulation, where the model can always tell them apart. Although the underlying reason remains uncertain, one plausible explanation for this behaviour is that the model’s ability to solve next-observation prediction tasks depends on their alignment with training sequences: for instance, it is plausible that Chameleon’s data rarely features two identical adjacent images.
3 Bootstrapping a World Model from a Dynamics Model in VLMs
@@
Since we showed in Section 2 that Chameleon-7B displays a higher proclivity as for action prediction than next-observation prediction, we first verify that this tendency is intensified when Chameleon-7B is fine-tuned on image editing trajectories (Section 3.1), as this results in the VLM acting reliably as a dynamics model. Motivated by this, we propose two strategies to leverage the VLMs as dynamics models to enhance VLMs as world models: (i) generating synthetic trajectories by annotating large-scale key-frame pairs from videos with actions predicted by the dynamics model, then using these as weak supervision to train the world model (Section 3.2); and (ii) using the dynamics model as a verifier at test time to score candidate next observations sampled from the world model (Section 3.3).
@@
Figure 3: Heatmap visualization of image token weights predicted by the recognition model on examples from UCF-101, Something-Something, MagicBrush, and Kubric.
3.1 Fine-tuning Chameleon as a Dynamics Model
@@
First, we fine-tune Chameleon as a Dynamics Model (CDM) $p_{\text{CDM}}(a\mid o_{s},o_{t})$, which predicts the probability of an action given the previous and next observations. As training data, we rely on high-quality triplets from Aurora (krojer2024aurora, ) and the action recognition track of EPIC-Kitchen (damen2018epic-kitchens, ), which is based on videos with an egocentric view. We use the first and last frame in the EPIC-Kitchen video clips as the source and target observation $o_{s}$ and $o_{t}$ and the annotated action as $a$. We provide full details on CDM training data and experimental setting Appendix A.8.1. Foreshadowing the results in Section 4.2, this significantly enhances action-prediction capabilities of Chameleon by a wide margin.
@@ 3.2 Weakly Supervised Learning from Unlabelled Videos
Taking advantage of the resulting high-quality CDM, we then explore the first of our strategies to bootstrap a world model in VLMs: we annotate pairs of motion key-frames of unlabelled videos with a textual description of the action with the CDM. To ensure both scale and quality, we collect approximately 45 hours of video from Moments-in-Time (monfortmoments-MIT, ), Kinetics-700 (kinetics700, ; carreira2019kinetics700Note, ), and UCF-101 (soomro2012ucf101, ), all of which consist of curated clips focused on human actions. To ensure the selected pairs of motion key-frames are meaningful, i.e., they express a valid action, we then calculate the optical flow to quantify the dynamics per frame in the video clips, and select the top-$K_{f}$ frames while ensuring that the interval between two selected frames is $I_{f}$. Specifically, we set $I_{f}=20$ and $K_{f}=6$ for all three datasets. This results in approximately 20K, 46K, and 21K annotated trajectory triplets from Moments-in-Time, Kinetics-700, and UCF-101, respectively. Finally, we apply a filtering strategy to further guarantee the quality of the resulting triplets. We use the CDM’s predicted likelihood for each trajectory triplet ($o_{s}$, $a_{\text{CDM}}$, $o_{t}$) as a score, and apply stratified Top-K sampling [^4] The details of this algorithm are provided in Appendix A.5. to select a subset of CDM-annotated trajectory triplets. We show statistics of the scores and action classes for the processed triplets in Figure 11. We also provide one example for each dataset in Figure 3.
Fine-tuning Chameleon as a World Model.
Afterwards, we fine-tune Chameleon as a World Model (CWM), $p_{\text{CWM}}(o_{t}\mid a,o_{s})$ on both Aurora’s supervised triplets $\mathcal{D}_{\text{sup}}$ and unsupervised triples $\mathcal{D}_{\text{unsup}}$ with actions sampled from the CDM. The world model CWM is trained with maximum likelihood estimation as an objective:
$\min_{\theta}\;\mathbb{E}_{(a,o_{s},o_{t})\sim\mathcal{D}_{\text{sup}}}\left[-\log p_{\theta}(o_{t}\mid a,o_{s})\right]+\,\mathbb{E}_{(o_{s},o_{t})\sim\mathcal{D}_{\text{unsup}}}\left[\mathbb{E}_{\hat{a}\sim p_{\text{CDM}}(a\mid o_{s},o_{t})}\left[-\log p_{\theta}(o_{t}\mid\hat{a},o_{s})\right]\right],$ (1)
where $\theta$ are the parameters for CWM, and $\hat{a}$ is action sampled from the CDM.
@@
Nevertheless, the objective in Equation 1 is limited by treating all regions of the target observation equally, even if some of them remain identical to the source whereas others change. This may result in degenerate solutions such as always copying the source. As an alternative, we therefore propose a novel training objective for world models that overcomes this assumption. This objective weights the loss of next-observation image tokens based on their importance. The intuition is that not all image patches in source and target observations contribute equally to modelling real-world transitions; instead, the model should focus on patches most indicative of the action’s consequences. To this end, we leverage a recognition model $f_{\text{rec}}(o_{s},o_{t})$, which outputs token-level weights aligned with Chameleon’s image token representations. These weights modulate the loss during training, emphasising learning on semantically meaningful regions and down-weighting irrelevant ones. We formulate our alternative objective as:
$\min_{\theta}\;\sum_{l=1}^{L}f_{\text{rec}}(o_{s},o_{t})^{(l)}\cdot\left(-\log p_{\theta}(o_{t}^{(l)}\mid o_{t}^{(<l)},o_{s},a)\right),$ (2)
where $\theta$ are the parameters of CWM and a $L$ is the number of tokens used to represent an image in Chameleon. $o_{t}^{(l)}$ and $o_{t}^{(<l)}$ represent the image tokens of $o_{t}$ at position $l$ and the history of previous positions, respectively. For simplicity, we use the pre-trained vector-quantised model of Chameleon as the recognition model, by computing the squared L2 norm of pre-quantized features $\mathbf{z}_{o_{s}}\in Z_{o_{s}}$ and $\mathbf{z}_{o_{t}}\in Z_{o_{t}}$ where $Z_{o_{s}}$ and $Z_{o_{t}}$ are the sets of features of source and target observations, respectively. We visualise the token weights in Figure 3, which capture the effects of acting on the source observation to yield the target one.

### v2

# Bootstrapping Action-Grounded Visual Dynamics in Unified Vision–Language Models

@@ Abstract
Can unified vision–language models (VLMs) perform forward dynamics prediction (FDP), i.e., predicting the future state (in image form) given the previous observation and an action (in language form)? We find that VLMs struggle to generate physically plausible transitions between frames from instructions. Nevertheless, we identify a crucial asymmetry in multimodal grounding: fine-tuning a VLM to learn inverse dynamics prediction (IDP)—effectively captioning the action between frames—is significantly easier than learning FDP. In turn, IDP can be used to bootstrap FDP through two main strategies: 1) weakly supervised learning from synthetic data and 2) inference time verification. Firstly, IDP can annotate actions for unlabelled pairs of video frame observations to expand the training data scale for FDP. Secondly, IDP can assign rewards to multiple samples of FDP to score them, effectively guiding search at inference time. We evaluate the FDP resulting from both strategies through the task of action-centric image editing on Aurora-Bench with two families of VLMs. Despite remaining general-purpose, our best model achieves a performance competitive with state-of-the-art image editing models, improving on them by a margin between $7\%$ and $13\%$ according to GPT4o-as-judge, and achieving the best average human evaluation across all subsets of Aurora-Bench. [^1] The code and models developed in this paper are available at https://github.com/yfqiu-nlp/vlm-world-model..
Bootstrapping Action-Grounded Visual Dynamics
in Unified Vision–Language Models
1Yifu Qiu, 3Yftah Ziser, 2Anna Korhonen, 1Shay B. Cohen, 1,2Edoardo M. Ponti 1Institute for Language, Cognition and Computation, University of Edinburgh 2Language Technology Lab, University of Cambridge 3NVIDIA {yifu.qiu,scohen,eponti}@ed.ac.uk
Figure 1: A high-level illustration of our two strategies to bootstrap Forward Dynamics Prediction from Inverse Dynamics Prediction in unified Vision–Language Models: (i) synthesising trajectories for weak supervision (left) and (ii) inference-time verification of candidate future observations (right).
@@ 1 Introduction
Figure 2: Percentage of times 9 VLMs assign higher probability to observation–action–observation Reference trajectories compared with 4 types of Negative (i.e., adversarially manipulated) trajectories, for both forward dynamics and inverse dynamics prediction. Higher values are better.
World models are instrumental in training embodied agents to endow them with specific abilities, such as planning and simulation (Qin et al., 2024; Brohan et al., 2023; Huang et al., 2022; Li et al., 2025; Reed et al., 2022a; Yang et al., 2023; Hafner et al., 2025, inter alia). However, learning a specialised world model is challenging due to the scarcity of real-world data (Liu et al., 2024; Motamed et al., 2025). Conversely, a promising alternative is endowing pre-existing unified [^2] Here ‘unified’ means models capable of interleaving text and images during generation architecturally. vision-language models (VLMs) with world modelling abilities. In fact, VLMs are already imbued with plenty of real-world knowledge of both action (in language form) and perception (in vision form), because of their large-scale pre-training.
@@
Firstly, we probe whether unified VLMs already contain reliable forward dynamics models (FDP), i.e., the ability to predict the next image frame given the previous image frame and an action expressed as a language instruction. For this assessment, we limit ourselves to single-step trajectories, as a first step towards longer horizons. Based on our evaluation, we empirically demonstrate that existing VLMs do not prefer ground-truth trajectories compared to adversarially generated ones. Hence, we verify that the world model implicit in the original VLMs per se is not well grounded on real-world forward dynamics (Gao et al., 2024; Qiu et al., 2024; Abdou et al., 2021).
@@
Surprisingly, we also find that predicting the action $a\in\mathcal{A}$ taking place between observations $o\in\mathcal{O}$—also known as inverse dynamics prediction (IDP; $\mathcal{O}_{t}\times\mathcal{O}_{t+1}\rightarrow\mathcal{A}_{t}$)—via supervised fine-tuning is substantially easier than FDP ($\mathcal{O}_{t}\times\mathcal{A}_{t}\rightarrow\mathcal{O}_{t+1}$). Inspired by this finding, we propose two strategies to bootstrap the FDP from the IDP in a unified VLM, namely (i) learning from synthetic trajectories in videos automatically labelled with actions by the dynamics model; and (ii) test-time verification of predicted observations sampled from the FDP through the IDP. Figure 1 illustrates our two strategies.
@@
For the weak supervision strategy, we use a VLM fine-tuned for IDP on the Aurora dataset Krojer et al. (2024) to annotate with linguistic actions ($\hat{a}\in\mathcal{A}$) motion key-frame pairs extracted from 45 hours of unlabelled real-world videos. These are sourced from movements-in-time (Monfort et al., 2019), Kinetics700 (Kay et al., 2017; Carreira et al., 2019) and UCF-101 (Soomro et al., 2012). Together with the ground-truth trajectories in Aurora, the synthesised trajectory triplets are then used for weakly supervised fine-tuning of the very same VLM for FDP. To encourage training to focus on image regions that change the most, we additionally propose a loss-weighting method which weights the loss of each image token according to the visual difference between the ground-truth previous and next observations, as estimated by a recognition model. Instead, in the verification strategy, we use the IDP to assign scores to multiple candidate samples generated by the FDP. Selecting the highest-score prediction can effectively guide search at inference time.
@@
We conduct a thorough evaluation of both strategies to bootstrap FDP from IDP on several datasets from Aurora-Bench (Krojer et al., 2024): MagicBrush, Action-Genome, Something-Something, WhatsUp and Kubric. We experiment with two families of unified VLMs for FDP, Chameleon-7B and Liquid-8B. Training on trajectories synthesised by IDP, our FDP can achieve an overall performance superior to state-of-the-art diffusion models specialised for image editing, both in terms of GPT4o-as-a-judge and human evaluation. Inference-time verification can also improve FDP to a comparable degree as trajectory synthesis, providing an effective training-free bootstrapping method.
While limited to single-step action–observation trajectories, this work offers promising early evidence that unified VLMs can be successfully endowed with FDP and hence may be suitably developed into long-horizon world models in the future.
@@
2 Unified VLMs Lack a Consistent Preference for Real-World Trajectories
@@
To understand whether off-the-shelf VLMs are already suitable for FDP, the first research question we investigate in this paper is: To what extent do VLMs exhibit a preference for sequences of actions and observations that align with real-world trajectories?
@@
To address this question, we evaluate 9 different VLMs on ground-truth trajectories from 5 subsets of Aurora-Bench (Krojer et al., 2024): MagicBrush, Something-Something, Action-Genome, Whatsup, and Kubric. Each subset contains 50 trajectory triplets of the form ($o_{s},a,o_{t}$), where $o_{s}$ is the source observation (image frame), $a$ the action (text), and $o_{t}$ the target observation. [^3] We choose these 9 VLMs using the following criteria: 1) they are publicly accessible and 2) they have been exposed to interleaved multimodal data during their pre-training.
@@
We then manually curate four types of negative trajectories using rule-based manipulations. The first kind is an action-level manipulation. 1) Random Action: for a given pair of observations, we substitute the original action with another randomly sampled within the same subset. We also perform three observation-level manipulations. 2) Random Observation: we randomly substitute the target observation with another in the same subset. 3) Copy Observation: we copy the source observation as the target observation. 4) Inverse Observation: we swap the source and target observations.
@@
In Figure 2, we compare the log-likelihood that VLMs assign to each ground-truth trajectory (Reference) against its corresponding manipulated one (Negatives). We evaluate the VLMs in two tasks: action prediction (i.e., as an inverse-dynamics model) and next-observation prediction (i.e., as a forward-dynamics model). For each kind of negative trajectory, we report the percentage of samples where the model favours the reference trajectory over the negative trajectory. From Figure 2, it emerges that VLMs display no clear preference for the ground-truth trajectories in a zero-shot setting (around $50\%$).
@@
In the action prediction task (right panel), there is a slightly higher tendency to favour the ground-truth over the group with random actions; however, even in the best case, Qwen2.5-VL-7B prefers the reference in only 60.08% of the samples. The only negative group that seems to be identifiable for VLMs is observation copying, where Qwen2.5-VL-7B has 67.34% of correct preference. In the next-observation prediction task (left panel), the VLM mostly fails in effectively differentiating the ground truth from the negatives. Although the underlying reason remains uncertain, one plausible explanation for this behaviour is that the model’s ability to solve next-observation prediction tasks depends on their alignment with training sequences: for instance, it is plausible that Chameleon’s data rarely features two identical adjacent images. We provide a discussion breaking down Chameleon’s performance in Appendix D.
3 Bootstrapping FDP from IDP
Overall, the results from Section 2 reveal that off-the-shelf VLMs are not suitable for FDP and IDP per se; however, they also show that IDP is a more feasible task, as VLMs already achieve accuracies above random chance in a zero-shot setting. Possibly, IDP abilities may be even improved with a small amount of fine-tuning. This underlies the key intuition behind our work: can we leverage this asymmetry to bootstrap forward-dynamics abilities from inverse-dynamics ones within a unified VLM?
@@ 3 Bootstrapping FDP from IDP
We propose two strategies to leverage VLMs fine-tuned for IDP to enhance their own FDP abilities: (i) generating synthetic trajectories by annotating large-scale key-frame pairs from videos with actions predicted by IDP, then using these as weak supervision to train for FDP (Section 3.2); and (ii) using the IDP as a verifier at test time to score candidate next observations sampled from the FDP (Section 3.3).
@@
Figure 3: Heatmap visualization of image token weights predicted by the recognition model on examples from AG, Something-Something, MagicBrush, and Kubric, and UCF-101, Kinetics700 and MIT.
3.1 Inverse-dynamics Modelling
@@
First, we fine-tune the unified VLM as an inverse-dynamics model (IDM) $p_{\text{IDM}}(a\mid o_{s},o_{t})$, which predicts the probability of an action given the previous and next observations. As training data, we rely on high-quality triplets from Aurora (Krojer et al., 2024) and the action recognition track of EPIC-Kitchen (Damen et al., 2018), which is based on videos with an egocentric view. We use the first and last frame in the EPIC-Kitchen video clips as the source and target observation $o_{s}$ and $o_{t}$ and the annotated action as $a$. We provide full details on IDM training data and setting in Appendix H.1.
@@ 3.2 Weakly Supervised Learning from Unlabelled Videos
Taking advantage of the resulting high-quality IDM, we then explore the first of our strategies to bootstrap the FDP in VLMs: we annotate pairs of motion key-frames of unlabelled videos with a textual description of the action with the IDM. To ensure both scale and quality, we collect approximately 45 hours of video from Moments-in-Time (Monfort et al., 2019), Kinetics-700 (Kay et al., 2017; Carreira et al., 2019), and UCF-101 (Soomro et al., 2012), all of which consist of curated clips focused on human actions. To ensure the selected pairs of motion key-frames are meaningful, i.e., they express a valid action, we then calculate the optical flow to quantify the dynamics per frame in the video clips, and select the top-$K_{f}$ frames while ensuring that the interval between two selected frames is $I_{f}$. Specifically, we set $I_{f}=20$ and $K_{f}=6$ for all three datasets. This results in approximately 20K, 46K, and 21K annotated trajectory triplets from Moments-in-Time, Kinetics-700, and UCF-101, respectively. Finally, we apply a filtering strategy to further guarantee the quality of the resulting triplets. Specifically, we apply stratified top-$k$ sampling based on the IDM’s predicted likelihood for each trajectory triplet ($o_{s}$, $\hat{a}_{\text{IDM}}$, $o_{t}$) [^4] The details of this algorithm are provided in Appendix E. to select a subset of triplets. We show statistics of the scores and action classes for the selected triplets in Figure 9. We also provide one example for each dataset in Figure 3.
FDP.
Finally, we fine-tune a VLM as a forward-dynamics model (FDM), $p_{\text{FDM}}(o_{t}\mid a,o_{s})$ on both Aurora’s supervised triplets $\mathcal{D}_{\text{sup}}$ and unsupervised triples $\mathcal{D}_{\text{unsup}}$ with actions sampled from the IDM. Normally, a FDM would be trained with maximum likelihood estimation as an objective:
$\min_{\theta}\;\mathbb{E}_{(a,o_{s},o_{t})\sim\mathcal{D}_{\text{sup}}}\left[-\log p_{\theta}(o_{t}\mid a,o_{s})\right]+$ (1)
$\mathbb{E}_{(o_{s},o_{t})\sim\mathcal{D}_{\text{unsup}}}\left[\mathbb{E}_{\hat{a}\sim p_{\text{IDM}}(a\mid o_{s},o_{t})}\left[-\log p_{\theta}(o_{t}\mid\hat{a},o_{s})\right]\right]$
where $\theta$ are the parameters for FDM, and $\hat{a}$ is an action sampled from the IDM.
@@
Nevertheless, the objective in Equation 1 is limited by treating all regions (i.e., image patches) of the target observation equally, even if some of them remain identical to the source whereas others change. This may result in degenerate solutions such as always copying the source. As an alternative, we therefore propose a novel training objective for FDM that overcomes this assumption. This objective weights the loss of next-observation image tokens based on their importance. The intuition is that not all image patches in source and target observations contribute equally to modelling real-world transitions; instead, the model should focus on patches most indicative of the action’s consequences. To this end, we leverage a recognition model $f_{\text{rec}}(w;o_{s},o_{t})$, which outputs token-level weights that represent the similarity between source and target patches. These weights modulate the loss during training, emphasising learning on semantically meaningful regions and down-weighting irrelevant ones. We formulate our objective as:
$\min_{\theta}\;\sum_{l=1}^{L}f_{\text{rec}}(w;o_{s},o_{t})^{(l)}\cdot\left(-\log p_{\theta}(o_{t}^{(l)}\mid o_{t}^{(<l)},o_{s},a)\right),$ (2)
where $\theta$ are the parameters of VLM as FDM and a $L$ is the number of image tokens. $o_{t}^{(l)}$ and $o_{t}^{(<l)}$ represent the image tokens of $o_{t}$ at position $l$ and the history of previous positions, respectively. For simplicity, we use the pre-trained vector-quantised encoder of the unified VLMs as the recognition model, by computing the squared L2 norm of pre-quantized features $\mathbf{z}_{o_{s}}\in Z_{o_{s}}$ and $\mathbf{z}_{o_{t}}\in Z_{o_{t}}$ where $Z_{o_{s}}$ and $Z_{o_{t}}$ are the sets of features of source and target observations, respectively. We visualise the token weights in Figure 3, which capture where acting on the source observation yields the largest effects on the target one.